OpenStudy (hoopcoach99):

Give reasons for each step of the deductive argument that if x – 5 = 2(x + 6), then x = -17. x – 5 = 2(x + 6) Original problem x – 5 = 2x + 12 1) -5 = x + 12 2) -17 = x 3) x = -17 4) Answer choices: Reflexive property Symmetric property Transitive property Addition property of equality Subtraction property of equality Multiplication property of equality Division property of equality Substitution property Distributive property

OpenStudy (anonymous):

hi! your answer is substitution property! I majored in mathematics, (:

OpenStudy (hoopcoach99):

Is that for the 1st step. I have to give one for each step?

OpenStudy (anonymous):

yes! (:

OpenStudy (studygurl14):

x - 5 = 2(x + 6) x - 5 = 2x + 12 (distriubutive prop)) -5 = x + 12 (subtraction prop of =, substitution) -17 = x (subtraction prop of =, substitution) x = -17 (symmetric property)

OpenStudy (hoopcoach99):

Thank you.

OpenStudy (anonymous):

wow..

OpenStudy (hoopcoach99):

So when you write the subtraction property of equality, I also write substitution. Can it be 2 different properties

OpenStudy (studygurl14):

it's both, because the step is composed of two parts

OpenStudy (hoopcoach99):

Ok, now I understand, This stuff is not easy.

OpenStudy (hoopcoach99):

Do you know how to Write the definition of the word acute angle as a biconditional statement.